OpenStudy (anonymous):

If Joe wanted to create a function that modeled a base of 8 and what exponents were needed to reach specific values, how would he set up his function?

OpenStudy (anonymous):

\[f(x)=x^8\] \[f(x)=\log_8x\] \[f(x)=\log_x8\] \[f(x)=8^x\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Help meee!!

zepdrix (zepdrix):

Closed? Figure it out? :3